Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

GMAT Prep Question

Expert replies
by [email protected] » Tue Sep 25, 2007 11:39 am
As there are total 4 letters and 4 envelopes
1 letter can be put in 4 diff envelopes in which only 1 way of putting is correct
So total no of ways letters can be put = 4*4= 16
As 1 letter can be put in the correct envelope in 1 way and there are four envelopes
the prob = 4/16 = 1/4
whats the OA
Join the discussion
Source: — Problem Solving |

by raulverde » Wed Sep 26, 2007 9:49 am
There are four Letters {L1,L2,L3,L4}
There are four Envelopes{E1,E2,E3,E4}

Ways to put 1 Letter in the correct Envelope ?

(L1 GOES INTO E1) OR (L2 GOES INTO E2) OR (L3 GOES INTO E3) OR (L4 GOES INTO E4)

(Event of picking the right combination ) = 1/4
(Event of picking the wrong combination) = 2/3 * 1/2 *1/1 = 1/3

Probability of all these events is the same 1/12 .

1/12 + 1/12 + 1/12 + 1/12 = 4*(1/12) = 1/3
Join the discussion

by wongee » Sat Sep 29, 2007 12:54 pm
hey I lost the method in the portion:

(Event of picking the right combination ) = 1/4
(Event of picking the wrong combination) = 2/3 * 1/2 *1/1 = 1/3 (how does this come about?)

Probability of all these events is the same 1/12 . (and this)

1/12 + 1/12 + 1/12 + 1/12 = 4*(1/12) = 1/3
(this i get)

THANKS IN ADVANCE!
Join the discussion

by raulverde » Mon Oct 01, 2007 11:51 pm
wongee,

Say you have 3 Envelopes left (e1,e2,e3) the corresponding letters for them are (l1,l2,l3)

now if we were to pick a wrong combinations . For the first wrong combination. we would have 3 choices of envelopes and 2 choices of letters ( We dont include the 1 right letter )

Hence 2/3

The next wrong combination would be 1/2 and the last one 1/1

HTH
Join the discussion

by ash g » Fri Mar 07, 2008 5:59 pm
hi, I would like to slightly twist the question in an effort to better understand probability.

[ Refering to raulverde explanation]
If the QUESTION WAS: What is the probablity that none of the letters match the correct envelope, would the solution be as simple as below or would it be more complicated.

(E1 doesnt match) & (E2 doesnt match) & (E3 doesnt match) & (E4 doesnt match)

(3/4) (2/3) (1/2) (1) = 1/4 <<==Answer
!E1 !E2 !E3 !E4

The reason for doubt is simply that it may so happen that L2 matched with E1, so the second term could be (3/3) rather than (2/3). Do we cater for such a event ?
Join the discussion

by ash g » Fri Mar 07, 2008 6:09 pm
hi, I would like to slightly twist the question in an effort to better understand probability.

[ Refering to raulverde explanation]
If the QUESTION WAS: What is the probablity that none of the letters match the correct envelope, would the solution be as simple as below or would it be more complicated.

(E1 doesnt match) & (E2 doesnt match) & (E3 doesnt match) & (E4 doesnt match)

(3/4) (2/3) (1/2) (1) = 1/4 <<==Answer
!E1 !E2 !E3 !E4

The reason for doubt is simply that it may so happen that L2 matched with E1, so the second term could be (3/3) rather than (2/3). Do we cater for such a event ?
Join the discussion

by GMATters1001 » Sat Dec 20, 2008 10:04 pm
anyone have a better explanation for this?
Join the discussion

by ronniecoleman » Sun Dec 21, 2008 7:59 am
Nice question!!!

Ok

Choose one letter which should be have a right address... 4C1

left three letters
Total ways: 3!
but it contains ways where one address is correct, so total 3 ways
+ where all letters are correct

so 3!-4

Now

( 4C1 *(3!-4) ) / 4!

= 1/ 3
Admission champion, Hauz khaz
011-27565856
Join the discussion

by gmatapril » Fri Mar 25, 2011 3:17 pm
why we have to do the last step 1/12* 4. since we are asked about probability of only one letter to be put in correct envelope address
raulverde wrote:There are four Letters {L1,L2,L3,L4}
There are four Envelopes{E1,E2,E3,E4}

Ways to put 1 Letter in the correct Envelope ?

(L1 GOES INTO E1) OR (L2 GOES INTO E2) OR (L3 GOES INTO E3) OR (L4 GOES INTO E4)

(Event of picking the right combination ) = 1/4
(Event of picking the wrong combination) = 2/3 * 1/2 *1/1 = 1/3

Probability of all these events is the same 1/12 .

1/12 + 1/12 + 1/12 + 1/12 = 4*(1/12) = 1/3
Join the discussion